# TRABSA: Interpretable Sentiment Analysis of Tweets using Attention-based BiLSTM and Twitter-RoBERTa

### v2

@@ Results
Word Embeddings Models Macro Average Precision Macro Average Recall Macro Average F1-score Accuracy
BoW RF 75% 61% 61% 67%
BoW NB 63% 62% 62% 64%
BoW SVM 76% 76% 76% 77%
BoW GBM 82% 79% 79% 81%
BoW LGBM 83% 81% 81% 83%
BoW XGBoost 77% 75% 75% 78%
BoW Catboost 80% 74% 75% 77%
BoW LGBM+KNN+MLP 82% 81% 81% 84%
BoW RF+KNN+MLP 74% 73% 74% 75%
BoW GBM+RF Stacking Classifier 78% 76% 76% 78%
BoW GBM+RF Voting Classifier 78% 69% 70% 72%
BoW Single Hidden Layer NN 71% 69% 69% 72%
BoW 3 Hidden Layers NN 68% 67% 67% 70%
TFIDF RF 69% 60% 60% 66%
TFIDF NB 65% 56% 56% 61%
TFIDF SVM 74% 70% 70% 73%
TFIDF GBM 80% 77% 77% 79%
TFIDF LGBM 80% 77% 78% 80%
TFIDF XGBoost 68% 67% 67% 70%
TFIDF Catboost 74% 72% 72% 74%
TFIDF LGBM+KNN+MLP 79% 77% 77% 79%
TFIDF RF Bagging 76% 48% 43% 56%
TFIDF RF+KNN+MLP 75% 71% 72% 74%
TFIDF GBM+RF Stacking Classifier 77% 76% 76% 78%
TFIDF GBM+RF Voting Classifier 75% 69% 70% 73%
TFIDF Single Hidden Layer NN 68% 67% 67% 69%
TFIDF 3 Hidden Layers NN 93% 92% 92% 93%
word2vec RF 51% 49% 49% 56%
word2vec NB 41% 34% 19% 36%
word2vec SVM 67% 49% 45% 57%
word2vec GBM 51% 50% 50% 56%
word2vec LGBM 53% 49% 47% 57%
word2vec XGBoost 55% 50% 48% 56%
word2vec Catboost 71% 49% 45% 57%
word2vec LGBM+KNN+MLP 53% 46% 41% 53%
word2vec RF+KNN+MLP 55% 49% 43% 57%
word2vec GBM+RF Stacking Classifier 46% 45% 45% 51%
word2vec GBM+RF Voting Classifier 47% 47% 47% 53%
word2vec Single Hidden Layer NN 36% 45% 40% 53%
word2vec 3 Hidden Layers NN 38% 48% 42% 56%
BERT RF 53% 51% 49% 58%
BERT NB 50% 51% 50% 53%
BERT SVM 52% 52% 52% 56%
BERT GBM 56% 54% 54% 60%
BERT LGBM 58% 57% 57% 61%
BERT XGBoost 56% 56% 56% 60%
BERT Catboost 56% 49% 46% 57%
BERT LGBM+KNN+MLP 61% 59% 59% 62%
BERT RF+KNN+MLP 60% 58% 57% 62%
BERT GBM+RF Stacking Classifier 55% 54% 54% 56%
BERT GBM+RF Voting Classifier 60% 58% 58% 66%
BERT Single Hidden Layer NN 61% 58% 59% 62%
BERT 3 Hidden Layers NN 62% 60% 60% 64%
SBERT RF 61% 48% 44% 54%
SBERT NB 53% 40% 32% 44%
SBERT SVM 56% 57% 56% 58%
SBERT GBM 55% 51% 49% 55%
SBERT LGBM 55% 52% 50% 56%
SBERT XGBoost 56% 56% 56% 57%
SBERT Catboost 69% 48% 42% 53%
SBERT LGBM+KNN+MLP 55% 49% 47% 53%
SBERT RF+KNN+MLP 56% 54% 54% 57%
SBERT GBM+RF Stacking Classifier 44% 44% 43% 48%
SBERT GBM+RF Voting Classifier 53% 50% 50% 54%
SBERT Single Hidden Layer NN 55% 55% 54% 59%
SBERT 3 Hidden Layers NN 56% 57% 57% 59%
RoBERTa RF 62% 62% 62% 64%
RoBERTa NB 63% 55% 52% 54%
RoBERTa SVM 65% 65% 65% 67%
RoBERTa GBM 63% 62% 63% 65%
RoBERTa LGBM 64% 63% 64% 66%
RoBERTa XGBoost 65% 63% 63% 66%
RoBERTa Catboost 63% 62% 61% 64%
RoBERTa LGBM+KNN+MLP 66% 66% 65% 66%
RoBERTa RF+KNN+MLP 58% 55% 55% 60%
RoBERTa GBM+RF Stacking Classifier 61% 61% 61% 63%
RoBERTa GBM+RF Voting Classifier 60% 60% 59% 60%
RoBERTa Single Hidden Layer NN 84% 84% 84% 84%
RoBERTa 3 Hidden Layers NN 84% 83% 83% 84%
RoBERTa BiLSTM+3 Hidden Layers NN 84% 84% 84% 85%
RoBERTa BilSTM+CNN 83% 81% 82% 83%
RoBERTa Proposed TRABSA Model 94% 93% 94% 94%
Compared to the best-performing traditional models like LightGBM (83%), stacked LGBM+KNN+MLP, and advanced hybrid state-of-the-art BiLSTM+3 Hidden Layers NN (85%), the TRABSA model achieves an impressive accuracy of 94%, indicating a substantial improvement of at least 11% over the closest competitors. The TRABSA model consistently outperforms others in accuracy and macro average precision, recall, and F1-score, demonstrating its robustness and effectiveness across different evaluation criteria. Even when compared to sophisticated neural network architectures like Single Hidden Layer NN (84%) and 3 Hidden Layers NN (84%), the TRABSA model exhibits a remarkable performance boost of 11.90%, further emphasizing its superiority in sentiment analysis tasks. The TRABSA model’s outstanding performance underscores its hybrid architecture’s efficacy, which integrates transformer-based mechanisms, attention mechanisms, and BiLSTM networks to capture nuanced sentiment patterns effectively. The findings suggest that while pre-trained transformer models like BERT and RoBERTa offer competitive performance, customized architectures like TRABSA tailored specifically for sentiment analysis tasks can yield substantial accuracy and predictive power improvements.
@@
Dataset Type Dataset Name DL Models Evaluation Metrics
@@
Extended Global COVID-19 Dataset Single Hidden Layer NN 97% 97% 97% 97%
3 Hidden Layers NN 97% 97% 97% 97%
BiLSTM+3 Hidden Layers NN 97% 97% 97% 97%
BiLSTM+CNN 11% 33% 17% 33%
Proposed TRABSA Model 98% 98% 98% 98%
Extended USA COVID-19 Dataset Single Hidden Layer NN 81% 81% 81% 83%
3 Hidden Layers NN 85% 83% 84% 85%
BiLSTM+3 Hidden Layers NN 85% 85% 85% 86%
BiLSTM+CNN 17% 33% 22% 51%
Proposed TRABSA Model 87% 86% 86% 87%
External Twitter Dataset Single Hidden Layer NN 93% 93% 93% 93%
3 Hidden Layers NN 92% 92% 92% 92%
BiLSTM+3 Hidden Layers NN 92% 92% 92% 92%
BiLSTM+CNN 53% 49% 46% 49%
Proposed TRABSA Model 97% 97% 97% 97%
External Reddit Dataset Single Hidden Layer NN 94% 93% 94% 94%
3 Hidden Layers NN 94% 94% 94% 94%
BiLSTM+3 Hidden Layers NN 94% 94% 94% 94%
BiLSTM+CNN 94% 94% 94% 94%
Proposed TRABSA Model 94% 93% 94% 95%
External Apple Dataset Single Hidden Layer NN 81% 82% 81% 84%
3 Hidden Layers NN 83% 81% 82% 85%
BiLSTM+3 Hidden Layers NN 87% 85% 86% 89%
BiLSTM+CNN 85% 83% 84% 87%
Proposed TRABSA Model 85% 86% 86% 90%
External US Airline Dataset Single Hidden Layer NN 93% 93% 93% 94%
3 Hidden Layers NN 94% 93% 93% 94%
BiLSTM+3 Hidden Layers NN 94% 93% 94% 94%
BiLSTM+CNN 94% 93% 93% 94%
Proposed TRABSA Model 95% 95% 95% 96%

### v3

@@ Results
Word Embedding Model Macro Average Precision Macro Average Recall Macro Average F1-score Accuracy Training Time (s) Inference Time (s)
BoW RF 75%$\pm$161%$\pm$261%$\pm$267%$\pm$16.73 0.03
BoW NB 63%$\pm$162%$\pm$262%$\pm$264%$\pm$10.82 0.01
BoW SVM 76%$\pm$276%$\pm$276%$\pm$277%$\pm$14.35 0.30
BoW GBM 82%$\pm$179%$\pm$179%$\pm$181%$\pm$1262.81 0.02
BoW LGBM 83%$\pm$181%$\pm$181%$\pm$183%$\pm$135.54 0.01
BoW XGBoost 77%$\pm$075%$\pm$175%$\pm$178%$\pm$086.50 0.02
BoW Catboost 80%$\pm$074%$\pm$175%$\pm$177%$\pm$1536.16 0.02
BoW LGBM+KNN+MLP 82%$\pm$181%$\pm$281%$\pm$284%$\pm$01422.53 45.43
BoW RF+KNN+MLP 74%$\pm$173%$\pm$274%$\pm$175%$\pm$11212.43 40.54
BoW GBM+RF Stacking Classifier 78%$\pm$076%$\pm$176%$\pm$178%$\pm$01364.06 50.57
BoW GBM+RF Voting Classifier 78%$\pm$069%$\pm$170%$\pm$172%$\pm$11250.37 48.30
BoW Single Hidden Layer NN 71%$\pm$469%$\pm$469%$\pm$472%$\pm$4646.34 24.53
BoW 3 Hidden Layers NN 68%$\pm$467%$\pm$467%$\pm$470%$\pm$4904.64 25.35
TFIDF RF 69%$\pm$160%$\pm$360%$\pm$366%$\pm$26.75 0.04
TFIDF NB 65%$\pm$156%$\pm$356%$\pm$361%$\pm$20.89 0.02
TFIDF SVM 74%$\pm$270%$\pm$370%$\pm$373%$\pm$25.15 0.46
TFIDF GBM 80%$\pm$177%$\pm$377%$\pm$379%$\pm$1273.51 0.05
TFIDF LGBM 80%$\pm$077%$\pm$378%$\pm$380%$\pm$136.12 0.02
TFIDF XGBoost 68%$\pm$167%$\pm$167%$\pm$170%$\pm$187.22 0.03
TFIDF Catboost 74%$\pm$172%$\pm$172%$\pm$174%$\pm$1542.44 0.03
TFIDF LGBM+KNN+MLP 79%$\pm$077%$\pm$177%$\pm$179%$\pm$01532.37 46.24
TFIDF RF Bagging 76%$\pm$048%$\pm$143%$\pm$156%$\pm$1764.34 35.34
TFIDF RF+KNN+MLP 75%$\pm$171%$\pm$372%$\pm$274%$\pm$01254.65 42.75
TFIDF GBM+RF Stacking Classifier 77%$\pm$076%$\pm$176%$\pm$178%$\pm$01352.53 52.65
TFIDF GBM+RF Voting Classifier 75%$\pm$069%$\pm$270%$\pm$173%$\pm$01283.23 50.75
TFIDF Single Hidden Layer NN 68%$\pm$367%$\pm$467%$\pm$469%$\pm$2650.34 22.43
TFIDF 3 Hidden Layers NN 93%$\pm$392%$\pm$492%$\pm$393%$\pm$4954.64 27.53
word2vec RF 51%$\pm$149%$\pm$249%$\pm$256%$\pm$16.82 0.05
word2vec NB 41%$\pm$134%$\pm$119%$\pm$136%$\pm$11.03 0.04
word2vec SVM 67%$\pm$049%$\pm$245%$\pm$357%$\pm$15.78 0.57
word2vec GBM 51%$\pm$150%$\pm$250%$\pm$256%$\pm$1282.10 0.08
word2vec LGBM 53%$\pm$149%$\pm$147%$\pm$157%$\pm$137.41 0.03
word2vec XGBoost 55%$\pm$150%$\pm$248%$\pm$256%$\pm$188.60 0.03
word2vec Catboost 71%$\pm$049%$\pm$145%$\pm$157%$\pm$1553.37 0.04
word2vec LGBM+KNN+MLP 53%$\pm$146%$\pm$241%$\pm$353%$\pm$11448.76 47.34
word2vec RF+KNN+MLP 55%$\pm$149%$\pm$343%$\pm$357%$\pm$11345.53 55.23
word2vec GBM+RF Stacking Classifier 46%$\pm$245%$\pm$245%$\pm$251%$\pm$11412.64 53.73
word2vec GBM+RF Voting Classifier 47%$\pm$147%$\pm$147%$\pm$153%$\pm$11350.23 28.78
word2vec Single Hidden Layer NN 36%$\pm$545%$\pm$440%$\pm$453%$\pm$3704.65 28.34
word2vec 3 Hidden Layers NN 38%$\pm$448%$\pm$442%$\pm$456%$\pm$31034.89 33.38
BERT RF 53%$\pm$251%$\pm$349%$\pm$358%$\pm$1805.43 45.53
BERT NB 50%$\pm$351%$\pm$250%$\pm$253%$\pm$11.12 0.07
BERT SVM 52%$\pm$252%$\pm$252%$\pm$256%$\pm$16.12 0.76
BERT GBM 56%$\pm$154%$\pm$254%$\pm$260%$\pm$2291.11 0.11
BERT LGBM 58%$\pm$257%$\pm$257%$\pm$261%$\pm$237.66 0.03
BERT XGBoost 56%$\pm$256%$\pm$256%$\pm$260%$\pm$289.10 0.04
BERT Catboost 56%$\pm$249%$\pm$246%$\pm$257%$\pm$2562.29 0.05
BERT LGBM+KNN+MLP 61%$\pm$159%$\pm$259%$\pm$262%$\pm$11623.65 65.34
BERT RF+KNN+MLP 60%$\pm$158%$\pm$257%$\pm$262%$\pm$11443.22 60.44
BERT GBM+RF Stacking Classifier 55%$\pm$254%$\pm$254%$\pm$256%$\pm$11523.75 70.23
BERT GBM+RF Voting Classifier 60%$\pm$158%$\pm$258%$\pm$266%$\pm$01452.45 67.85
BERT Single Hidden Layer NN 61%$\pm$358%$\pm$459%$\pm$462%$\pm$2945.53 35.64
BERT 3 Hidden Layers NN 62%$\pm$460%$\pm$460%$\pm$464%$\pm$41305.39 45.49
SBERT RF 61%$\pm$048%$\pm$344%$\pm$454%$\pm$26.95 0.06
SBERT NB 53%$\pm$240%$\pm$332%$\pm$444%$\pm$21.14 0.09
SBERT SVM 56%$\pm$157%$\pm$256%$\pm$258%$\pm$16.28 0.81
SBERT GBM 55%$\pm$251%$\pm$249%$\pm$255%$\pm$2302.43 0.15
SBERT LGBM 55%$\pm$252%$\pm$250%$\pm$256%$\pm$238.18 0.04
SBERT XGBoost 56%$\pm$256%$\pm$256%$\pm$257%$\pm$290.53 0.06
SBERT Catboost 69%$\pm$048%$\pm$342%$\pm$453%$\pm$2577.54 0.05
SBERT LGBM+KNN+MLP 55%$\pm$249%$\pm$247%$\pm$253%$\pm$21734.23 68.64
SBERT RF+KNN+MLP 56%$\pm$154%$\pm$254%$\pm$257%$\pm$11522.42 63.43
SBERT GBM+RF Stacking Classifier 44%$\pm$244%$\pm$243%$\pm$248%$\pm$11623.86 72.57
SBERT GBM+RF Voting Classifier 53%$\pm$250%$\pm$250%$\pm$254%$\pm$21553.54 74.67
SBERT Single Hidden Layer NN 55%$\pm$355%$\pm$454%$\pm$359%$\pm$3954.64 38.48
SBERT 3 Hidden Layers NN 56%$\pm$457%$\pm$357%$\pm$359%$\pm$21402.54 48.48
RoBERTa RF 62%$\pm$262%$\pm$262%$\pm$264%$\pm$17.23 0.08
RoBERTa NB 63%$\pm$155%$\pm$252%$\pm$354%$\pm$21.16 1.05
RoBERTa SVM 65%$\pm$165%$\pm$165%$\pm$167%$\pm$06.76 0.88
RoBERTa GBM 63%$\pm$262%$\pm$363%$\pm$265%$\pm$2314.44 0.17
RoBERTa LGBM 64%$\pm$263%$\pm$364%$\pm$266%$\pm$138.89 0.05
RoBERTa XGBoost 65%$\pm$163%$\pm$263%$\pm$266%$\pm$191.14 0.07
RoBERTa Catboost 63%$\pm$262%$\pm$361%$\pm$464%$\pm$2583.28 0.07
RoBERTa LGBM+KNN+MLP 66%$\pm$266%$\pm$265%$\pm$366%$\pm$21823.93 72.23
RoBERTa RF+KNN+MLP 58%$\pm$355%$\pm$255%$\pm$260%$\pm$11654.54 67.76
RoBERTa GBM+RF Stacking Classifier 61%$\pm$361%$\pm$361%$\pm$363%$\pm$21705.36 75.96
RoBERTa GBM+RF Voting Classifier 60%$\pm$260%$\pm$259%$\pm$360%$\pm$21653.78 73.47
RoBERTa Single Hidden Layer NN 84%$\pm$484%$\pm$484%$\pm$484%$\pm$41349.46 154.39
RoBERTa 3 Hidden Layers NN 84%$\pm$383%$\pm$483%$\pm$484%$\pm$31898.05 153.64
RoBERTa BiLSTM+3 Hidden Layers NN 84%$\pm$384%$\pm$384%$\pm$385%$\pm$23404.54 148.43
RoBERTa BilSTM+CNN 83%$\pm$281%$\pm$482%$\pm$383%$\pm$25328.73 178.64
RoBERTa Proposed TRABSA Model 94%$\pm$193%$\pm$294%$\pm$194%$\pm$13675.21 147.14
Table 4: Ablation test of the proposed TRABSA model
Word Embedding Model Macro Average Precision Macro Average Recall Macro Average F1-score Accuracy Training Time (s) Inference Time (s)
RoBERTa w/o111”w/o” stands for ”without,” indicating the absence of the specific component in the TRABSA model configuration. BiLSTM + Attention + 3 Dense Layers 76%$\pm$275%$\pm$375%$\pm$276%$\pm$31290.53 115.68
RoBERTa w/o BiLSTM + Attention Layer 81%$\pm$180%$\pm$180%$\pm$180%$\pm$23500.23 169.34
RoBERTa w/o Attention Layer 85%$\pm$284%$\pm$384%$\pm$284%$\pm$22250.15 145.85
RoBERTa w/o 3 Dense Layers 82%$\pm$281%$\pm$381%$\pm$382%$\pm$32245.64 150.14
RoBERTa w/o Dropout Layer 84%$\pm$184%$\pm$184%$\pm$184%$\pm$22200.70 140.45
RoBERTa LSTM instead of BiLSTM Layer 83%$\pm$282%$\pm$382%$\pm$283%$\pm$32380.90 142.30
Compared to the best-performing traditional models like GBM (81%), LightGBM (83%), stacked LGBM+KNN+MLP (84%), and advanced hybrid state-of-the-art BiLSTM+3 Hidden Layers NN (85%), the TRABSA model achieves an impressive accuracy of 94%, indicating a substantial improvement of at least 9% over the closest competitors. The TRABSA model consistently outperforms others in accuracy and macro average precision, recall, and F1-score, demonstrating its robustness and effectiveness across different evaluation criteria. Even when compared to sophisticated neural network architectures like Single Hidden Layer NN (84%) and 3 Hidden Layers NN (84%), the TRABSA model exhibits a remarkable performance boost of $\approx 10\%$, further emphasizing its superiority in SA tasks. The TRABSA model’s outstanding performance underscores its hybrid architecture’s efficacy, which integrates transformer-based mechanisms, attention mechanisms, and BiLSTM networks to capture nuanced sentiment patterns effectively. The findings suggest that while pre-trained transformer models like BERT and RoBERTa offer competitive performance, customized architectures like TRABSA tailored specifically for SA tasks can yield substantial accuracy and predictive power improvements.
The ablation study presented in Table LABEL:tab:ablation offers a detailed analysis of the proposed TRABSA model by examining the effects of removing or altering key components. The full model, not shown in this table, achieves the highest performance, with macro average precision and F1-score reaching 94%, macro average recall at 93%, and accuracy at 94%. This highlights the effectiveness of the BiLSTM, Attention, and Dense layers in capturing complex patterns from the data.
When the BiLSTM, Attention, and Dense layers are all removed, the model’s performance drastically drops to 76% for precision, 75% for recall, 75% for F1-score, and 76% for accuracy. The reduction in training time ($\approx$1290 seconds) and inference time ($\approx$116 seconds) indicates the computational savings from eliminating these layers, but the performance decline reveals their critical role in generalizing well to the data.
In the model where only the Attention layer is removed, the performance sees a significant improvement compared to the full removal of BiLSTM and Dense layers. This configuration achieves an 85% macro average precision, 84% recall, and 84% F1-score, with an accuracy of 84%. The training time is $\approx$2250 seconds, and the inference time is $\approx$146 seconds, suggesting that while the Attention layer adds value, its absence doesn’t drastically impair the model’s ability to capture temporal dependencies, likely due to the strength of the BiLSTM.
Interestingly, replacing the BiLSTM with a unidirectional LSTM results in a slight drop in performance across all metrics, with precision at 83%, recall at 82%, F1-score at 82%, and accuracy at 83%. As expected, the training and inference times ($\approx$2381 seconds and $\approx$142 seconds, respectively) show that unidirectional LSTMs are slightly more efficient but at the cost of losing the bidirectional context offered by the BiLSTM.
Moreover, the model without the Dropout layer exhibits similar performance to the attention-removed configuration, maintaining 84% across all metrics, but with a slightly faster training time ($\approx$2201 seconds) and inference time ($\approx$140 seconds). This suggests that Dropout contributes to regularization, but its absence does not significantly degrade performance, possibly due to the robustness of the RoBERTa embeddings and other layers.
The study underscores the importance of BiLSTM and Attention layers for optimal performance while also demonstrating the computational costs associated with these enhancements. The model without these components, while more computationally efficient, sacrifices accuracy, confirming the balance between complexity and performance in the proposed TRABSA model.
@@
Dataset Type Dataset Name DL Models Evaluation Metrics Training Time (s) Inference Time (s)
@@
Extended Global COVID-19 Dataset Single Hidden Layer NN 97%$\pm$097%$\pm$097%$\pm$097%$\pm$17365 491
3 Hidden Layers NN 97%$\pm$097%$\pm$097%$\pm$097%$\pm$17755 517
BiLSTM+3 Hidden Layers NN 97%$\pm$197%$\pm$197%$\pm$197%$\pm$15709 518
BiLSTM+CNN 11%$\pm$533%$\pm$417%$\pm$533%$\pm$48789 536
Proposed TRABSA Model 98%$\pm$098%$\pm$098%$\pm$098%$\pm$18288 518
Extended USA COVID-19 Dataset Single Hidden Layer NN 81%$\pm$381%$\pm$381%$\pm$383%$\pm$3870 58
3 Hidden Layers NN 85%$\pm$383%$\pm$184%$\pm$285%$\pm$3696 58
BiLSTM+3 Hidden Layers NN 85%$\pm$185%$\pm$185%$\pm$186%$\pm$01218 59
BiLSTM+CNN 17%$\pm$533%$\pm$522%$\pm$451%$\pm$1413 59
Proposed TRABSA Model 87%$\pm$186%$\pm$186%$\pm$187%$\pm$11081 47
External Twitter Dataset Single Hidden Layer NN 93%$\pm$393%$\pm$393%$\pm$393%$\pm$39891 495
3 Hidden Layers NN 92%$\pm$192%$\pm$192%$\pm$192%$\pm$14608 288
BiLSTM+3 Hidden Layers NN 92%$\pm$392%$\pm$392%$\pm$392%$\pm$38700 291
BiLSTM+CNN 53%$\pm$249%$\pm$446%$\pm$349%$\pm$41752 292
Proposed TRABSA Model 97%$\pm$197%$\pm$197%$\pm$197%$\pm$16602 287
External Reddit Dataset Single Hidden Layer NN 94%$\pm$393%$\pm$394%$\pm$394%$\pm$31494 90
3 Hidden Layers NN 94%$\pm$194%$\pm$194%$\pm$194%$\pm$22415 119
BiLSTM+3 Hidden Layers NN 94%$\pm$194%$\pm$294%$\pm$294%$\pm$12464 101
BiLSTM+CNN 94%$\pm$194%$\pm$094%$\pm$094%$\pm$01944 119
Proposed TRABSA Model 94%$\pm$193%$\pm$094%$\pm$095%$\pm$12200 94
External Apple Dataset Single Hidden Layer NN 81%$\pm$182%$\pm$281%$\pm$184%$\pm$396 11
3 Hidden Layers NN 83%$\pm$281%$\pm$382%$\pm$385%$\pm$155 10
BiLSTM+3 Hidden Layers NN 87%$\pm$285%$\pm$486%$\pm$389%$\pm$0140 12
BiLSTM+CNN 85%$\pm$183%$\pm$384%$\pm$287%$\pm$0130 11
Proposed TRABSA Model 88%$\pm$186%$\pm$286%$\pm$290%$\pm$0210 12
External US Airline Dataset Single Hidden Layer NN 93%$\pm$393%$\pm$393%$\pm$394%$\pm$21201 48
3 Hidden Layers NN 94%$\pm$293%$\pm$393%$\pm$394%$\pm$21166 53
BiLSTM+3 Hidden Layers NN 94%$\pm$193%$\pm$294%$\pm$194%$\pm$11012 46
BiLSTM+CNN 94%$\pm$393%$\pm$493%$\pm$494%$\pm$3842 40
Proposed TRABSA Model 95%$\pm$095%$\pm$095%$\pm$096%$\pm$1897 39
The robustness and generalizability of our TRABSA model are evident through its superior performance compared to a wide range of state-of-the-art models used in multiclass sentiment analysis (SA) on Twitter data. Table 6 summarizes these comparisons, showcasing how TRABSA consistently outperforms models across multiple datasets. Notably, on the global COVID-19 dataset, TRABSA achieves an exceptional macro average precision, recall, F1-score, and accuracy of 98%, significantly surpassing models such as Jlifi et al.[58], which utilized the Ens-RF-BERT approach and achieved a macro average F1-score of 94.03% and accuracy of 93.01%. Similarly, the model by Sazan et al.[57], which employed RoBERTa+fastText, attained an F1-score of 92.05% but still falls short when compared to TRABSA’s 95% F1-score on the US Airline dataset. Furthermore, the CNN-LSTM model proposed by Mohbey et al.[56] achieved 91.24% F1-score, showcasing a respectable result, but it is outperformed by TRABSA’s 98% on the global COVID-19 dataset.
Table 6: Summary of the proposed models in the state-of-the-art tweet sentiment analysis literature
Study Model Used Dataset Macro Average Precision Macro Average Recall Macro Average F1-score Accuracy
Qi & Shabrina (2023)[16] BoW+SVC UK COVID-19 Twitter Dataset 69.66% 70.33% 69.66% 71.00%
Ours TRABSA UK COVID-19 Twitter Dataset 94.00% 93.00% 94.00% 94.00%
dos Santos Neto et al., (2024)[49] BERT TripAdvisor 87.70% 88.20% 87.90% 88.20%
Brum & Volpe Nunes (2018)[50] BERT TweetSentBR 73.27% 72.75% 72.96% 72.75%
De Souza et al. (2018)[51] MultiFiT-Twitter LM Twitter NPS 72.43% 72.46% 72.43% 72.46%
Pilar et al. (2023)[52] Neighbor-sentiment InterTASS 57.76% 51.39% 54.39% 61.35%
Su & Kabala (2023)[53] GloVe100+LSTM 500k ChatGPT-related Tweets Jan-Mar 2023 81.10% 81.10% 81.10% 81.10%
Memiş et al. (2024)[54] Multiclass CNN model with pre-trained word embedding Turkish Financial Tweets - - - 72.73%
Kp et al. (2024)[55] Ensemble classifier Twitter API Dataset 91.29% 89.65% 87.32% 93.42%
Mohbey et al. (2024)[56] CNN-LSTM Monkeypox Tweets 91.24% 91.24% 91.24% 91.24%
Sazan et al., 2024)[57] RoBERTa+fastText US Airline Dataset 92.08% 92.02% 92.05% 92.02%
Ours TRABSA US Airline Dataset 95.00% 95.00% 95.00% 96.00%
Jlifi et al. (2024)[58] Ens-RF-BERT Hashtag Covid19 Tweets 94.03% 93.05% 94.03% 93.01%
Bhardwaj et al. (2024)[59] BoW+LR COV19Tweets Dataset 82.00% 81.80% 81.60% 81.80%
Ours TRABSA Global COVID-19 Dataset 98.00% 98.00% 98.00% 98.00%
What sets TRABSA apart is its consistent performance across different datasets, including both domain-specific (e.g., the US Airline dataset, where it achieved 96% accuracy) and global datasets. In contrast, other models often exhibit variability in performance depending on the dataset or sentiment categories. This ability to generalize across diverse contexts, such as pandemic-related tweets and US airline sentiment, highlights TRABSA’s robustness in handling complex multiclass SA tasks. The models compared in this table span various techniques, from traditional models such as BoW+SVC[16] to more modern DL architectures like CNN-LSTM[56] and transformer-based approaches such as BERT[49], yet none achieve the same level of performance as TRABSA.
Statistical Validation
To assess the performance of our proposed TRABSA model against the other top-performing models benchmarked in this study, we performed a 10-fold cross-validated two-tailed paired t-test, each with 50 epochs, on key evaluation metrics: accuracy, macro average precision, recall, and F1-score. By top-performing model, we refer to the best combination of word embedding and the model obtained from Table LABEL:tab3. Our null hypothesis ($H_{0}$) stated that there is no significant difference between the performance of each model and the TRABSA model for the respective metrics. In contrast, the alternative hypothesis ($H_{1}$) posited that a significant difference does exist. We utilized a significance level of $\alpha=5\%$, with a Bonferroni correction to account for multiple comparisons. The t-test results (see Table 7) revealed that for all metrics—accuracy, precision, recall, and F1-score—the TRABSA model showed statistically significant improvements compared to the other models ($p-values<0.05$ after adjustment) by rejecting the $H_{0}$. For example, the accuracy of the TRABSA model was significantly higher than that of BoW+RF, with a t-value of 108.7332 and a p-value of $2.39\times 10^{-15}$, suggesting a meaningful enhancement in performance. Similarly, the TRABSA model consistently demonstrated superior results with significant t-statistics and p-values for precision and recall. These findings robustly support the efficacy of the TRABSA model in delivering enhanced performance metrics compared to traditional models.
Table 7: 10-fold cross-validated paired t-tests comparing macro-average precision, recall, F1 scores, and accuracy of the top-performing models against the TRABSA model
Model t-value (Accuracy) p-value (Accuracy) t-value (Precision) p-value (Precision) t-value (Recall) p-value (Recall) t-value (F1 score) p-value (F1 score)
BoW_RF 108.7332 2.39E-15 59.19424 5.65E-13 135.5109 3.30E-16 128.4593 5.33E-16
BoW_NB 161.109 6.95E-17 95.1292 7.95E-15 120.9821 9.15E-16 105.8549 3.04E-15
BoW_SVM 85.36398 2.10E-14 67.73762 1.68E-13 66.34988 2.03E-13 51.12919 2.10E-12
BoW_GBM 54.78734 1.13E-12 35.68181 5.28E-11 48.87382 3.15E-12 52.17979 1.75E-12
BoW_LGBM 61.19279 4.19E-13 29.05617 3.30E-10 41.47385 1.37E-11 49.3594 2.88E-12
BoW_XGBoost 91.83215 1.09E-14 62.80512 3.32E-13 101.6696 4.37E-15 96.5374 6.96E-15
BoW_Catboost 82.98394 2.71E-14 42.12589 1.19E-11 120.3859 9.56E-16 63.42425 3.04E-13
BoW_LGBM+KNN+MLP 56.52622 8.54E-13 39.37306 2.19E-11 53.2787 1.45E-12 44.7578 6.94E-12
BoW_RF+KNN+MLP 115.5203 1.39E-15 54.42813 1.20E-12 83.80527 2.48E-14 81.01391 3.37E-14
BoW_GBM+RF Stacking Classifier 95.87232 7.41E-15 76.63942 5.55E-14 125.9309 6.38E-16 104.5914 3.39E-15
BoW_GBM+RF Voting Classifier 126.3773 6.18E-16 67.23083 1.80E-13 89.36265 1.39E-14 74.88783 6.83E-14
RoBERTa_Single Hidden Layer NN 46.94341 4.52E-12 31.96107 1.41E-10 41.38057 1.40E-11 30.30619 2.27E-10
RoBERTa_3 Hidden Layers NN 51.66872 1.91E-12 33.33187 9.70E-11 97.90134 6.14E-15 39.84824 1.96E-11
RoBERTa_BiLSTM+3 Hidden Layers NN 43.79562 8.43E-12 26.53577 7.41E-10 56.63986 8.39E-13 46.28162 5.14E-12
RoBERTa_BiLSTM+CNN 57.76621 7.03E-13 40.85393 1.57E-11 45.50036 5.98E-12 60.10854 4.92E-13